addition Property of equality: I'm not sure of the correct answer
x+y=13
22x + 36y = 356

Answer by checkley75(3666)

X+Y=13, Y=13-X
22X+36Y=356 NOW SUBSTITUTE (13-X) FOR Y IN THIS EQUATION & SOLVE FOR X
22X+36(13-X)=356
22X+468-36X=356
-14X=356-468
-14X=-112
X=-112/-14
X=8 ANSWER.
8+Y=13
Y=13-8
Y=5 ANSWER.
PROOF
22*8+36*5=356
176+180=356
356=356